# MAViS: Modular Autonomous Virtualization System for Two-Dimensional Semiconductor Quantum Dot Arrays

### v1

@@ II.3 Image processing
To extract from the experimentally acquired 2D CSDs information useful for automation we use an ensemble of five convolutional neural network (CNN) pixel classifiers [54] trained to distinguish each pixel as belonging to five different classes: horizontal, vertical, or diagonal transitions (corresponding to QD formed under left, right, or simultaneously under both plunger gates), interdots, and points where no transitions are present [45]. The pixel classifiers, trained using exclusively simulated data [55, 56], take as an input a small 2D plunger-plunger voltage scan obtained using a charge sensor, as shown in Fig. 1(c). It then labels each pixel within the scan with a 5D vector indicating the probability of belonging to each of the five possible charge transition classes.
Prior to running the pixel classifiers, the measured data is preprocessed. This involves a series of steps. First, the data is convolved with a Gaussian filter, and the gradient is taken. The resulting image is then mean normalized, and the normalized gradient is cropped to produce a $32\times 32$ pixel subimage.
Figure 2: Plunger gates orthogonalization and normalization. (a) Exemplary charge stability diagram spanned by the sensor-virtualized gates $\mathrm{vP}_{7}$ and $\mathrm{vP}_{3}$. In red, we overlay the output of the pixel classifier module for vertical transitions. The region circled in green indicates a single transition segment, while the blue dashed line indicates the composite line averaged over the three segments of the D7 transition line on the left side of the image. The red line in the center of the image has been erroneously marked by the classifier. (b) The Hough transform of the ML model output. The individual transitions become thin bands, which overlap at a single point corresponding to the composite line (marked in blue). The three additional peaks (marked in green) correspond to the individual transitions on the left side. Each pixel has a width of $0.17$ $\mathrm{mV}$ and a height of $0.35$ $\mathrm{mV}$. (c) The sum of the squares of the Hough transform, forming a bimodal distribution corresponding to the composite and individual transitions. (d) Inverse of the virtual matrix $M_{2}$ obtained from the Hough transforms (left) and a 2D CSD acquired with orthogonalized virtual plunger gates (right). (e) Inverse of the normalized virtual plunger matrix $M_{3}$ (left) and 2D CSD acquired with normalized virtual plunger gates. In the space spanned by $\mathrm{N}_{7}$ and $\mathrm{N}_{3}$, the charging voltages for the specific charge state are constant (20 $\mathrm{mV}$), allowing for uniform $x$- and $y$-axes.
@@
To identify the coordinates of the interdots, we implement a dynamic thresholding algorithm that selects the interdots based on the probabilities $\mathrm{P}(\ell)$ given by the ML module. The threshold $\delta$ is determined based on the mean $(\mu)$ and variance $(\sigma^{2})$ of a Gaussian fit to the log-transformed interdot class probability distribution,
@@
where $\eta$ is set dynamically in the range of $(2.5,4)$ in order to identify at least the expected number of interdots, see Fig. 1(f). In the end, a pixel is classified as belonging to an interdot if $\mathrm{P}(\ell)>\delta$. The final interdot coordinates in the 2D CSD are determined based on the center of mass of each cluster of points classified as an interdot, as illustrated in Fig. 1(g).
@@ II.4 Plunger virtualization
From a 2D CSD, the ML module identifies regions that correspond to horizontal and vertical transitions. For example, in Fig. 2(a), all pixels identified as belonging to vertical transitions are marked in red. This includes three left-side transitions and three right-side transitions, as well as an erroneously identified line in the center of the honeycomb. The ML output is then processed using the Hough transform, resulting in a representation where each transition becomes a thin band in the Hough space, with $\theta$ and $d$ corresponding to the angle and distance from the origin of the transition; see Fig. 2(b).
@@
The 2D CSDs often contain several nearly colinear transition lines that have slight offset relative to each other due to the presence of interdot transition, as shown in Fig. 2(a). This effect manifests itself visually in the Hough transformation as several small peaks (corresponding to each of the individual transitions) which merge together to form a single large peak [corresponding to the composite line that passes through the center of each of the nearly collinear transitions a dashed blue line in Fig. 2(a)], as shown in Fig. 2(b) with green and blue ovals, respectively. The ability to distinguish between these two cases is essential, as the slope of the composite line will be different from the slopes of the individual transitions themselves, and this difference will be magnified as the size of the interdots increases.
When this occurs, $H_{\text{sq}}(\theta)$ will have a bimodal distribution, with the peak closest to $\theta=\pi/2$ ($\theta=0$) for the case of horizontal (vertical) transitions corresponding to the composite line, and the other peak reflecting the slope of the individual transition segments, as shown in Fig. 2 (c). Our implementation of MAViS is designed to prefer the peak corresponding to individual transitions so that the horizontal and vertical transitions are orthogonal in the final virtualized space, although in principle the composite line can be chosen instead. After finding the angle $\theta$ corresponding to the individual transitions, the slope of these transitions is then obtained via basic trigonometric functions. This process is repeated for all nearest-neighbor and next-nearest-neighbor plunger-plunger pairs, yielding the elements of the matrix $M_{2}^{-1}$ in Table 1.
@@ II.5 Barrier virtualization
The next layer of MAViS is the virtualization of the barrier gates with respect to the QD levels. Because of the barrier-to-QD capacitive couplings, non-virtualized barriers result in a shift of the charge state within a CSD as the barrier gate is modified. Virtualized barriers are built to ensure that such a charge state, enclosed by four interdots, remains at the charge symmetry point despite varying the tunnel coupling between neighboring QDs.
Barrier virtualization involves determining the virtualization coefficients in Eq. (4) and Eq. (5) that will correct for the shifts of the interdot positions as the barrier gates are adjusted. The same ML models trained on 2D CSDs obtained from plunger-plunger sweeps are leveraged for this task. First, the interdot locations in the 2D CSD are determined as described in Sec. II.3. To mitigate errors in tracking individual interdots, the center of the honeycomb formed by a set of four interdots is tracked instead. If the pixel classifier fails to identify one of the tracked interdots in one of the scans, coordinates for the corresponding interdot from the previous scan are used instead. Then, we derive the rate at which the charge state shifts as the barrier voltage is changed using one of two methods.
@@
However, local tracking is only appropriate for finding the linear fit coefficient. For the quadratic fit observed in the ON regime, a global tracking method is implemented. The global method involves direct tracking of the trajectories of the interdot transition points as the barrier gate is pulsed over multiple scans. The center coordinates as a function of the corresponding barrier voltages are then fit to a linear or quadratic curve, as desired. While missing or misclassified points can negatively affect the quality of the fit, virtualization in the OFF regime ensures less drastic shifts in the ON regime, making the fit less prone to errors. Moreover, with a limited number of points involved in the fit, a single misidentified point can substantially affect the resulting fit and lead to a miscalibration, which makes this approach less robust with noisy data. To overcome this limitation and ensure a high-quality fit, the virtualization in the ON regime is established based on a significantly denser sampling of 2D CSD for every pair of gates. Since only the neighboring gates need to be virtualized at this step, a larger number of measurements per gate pair can be afforded compared to the OFF regime virtualization.
@@ III Results
MAViS is designed to autonomously virtualize a set of sensors, plungers, and barrier gates used to define QDs. It operates through five distinct modules, each corresponding to a specific virtualization layer outlined in Table 1. At its core is an ML pixel classifier that extracts high-level feature representations from experimental data. MAViS leverages horizontal and vertical transition classes to determine the plunger-plunger virtualization coefficients, while the interdot class is used to identify and track interdot transitions for barrier virtualization. The performance of each virtualization layer is discussed in the remainder of this section, while we focus on the scalability and time requirements of MAViS in the Supplemental Material [60].
@@ III.1 Charge sensors compensation
When searching for clear honeycomb-like patterns in 2D CSDs measured via charge sensing, it is critical to maintain the rf-reflectometry charge sensors tuned to their maximum-sensitivity point, i.e., at the flank of one of their Coulomb peaks, throughout the whole voltage scan. The high-sensitivity of charge sensing is ensured by calibrating the relative lever arm of each gate controlling the QD array to the charge sensors plunger gates. This step does not entail using ML algorithms and exploits only traditional analysis routines to track the charge sensors’ Coulomb peak position as a function of each barrier and plunger gate.
The charge sensor compensation matrix, $M_{1}$, is determined in the first virtualization layer by obtaining the crosstalk of each gate to the four charge sensors. This is achieved by extracting the slope of the Coulomb peak position as a function of each gate. The slope of the linear shift in the $\mathbf{S}$ vs. $\mathbf{P}$, and $\mathbf{S}$ vs. $\mathbf{B}$ space is extracted and fed into the matrix $M^{-1}_{1}$, as shown in the Supplemental Material [60]. As we move on to the next virtualization layers, only the sensor-virtualized $\mathbf{vP}$ and $\mathbf{vB}$ parameters are considered.
@@ III.2 Plunger gates virtualization
The plunger gate orthogonalization matrix $M_{2}$ is determined using the slope extraction method described in Sec. II.4. First, a 2D CSD acquired for each pair of plunger gates ($\mathrm{vP}_{i}$, $\mathrm{vP}_{j}$), with $i,j\in[1,\dots,10]$, and $i\neq j$, is processed by the ML module to detect the horizontal and vertical charge transition lines, see Fig. 2. Then, a Hough transform method is used to extract the slopes of all relevant charge transitions, see Fig. 2(b). The inverses of the slopes define the elements of the orthogonalization matrix; see the right panel in Fig. 2(d). When scanning the newly defined virtual plunger gates ($\mathrm{O}_{i}$, $\mathrm{O}_{j}$), with $i,j\in[1,\dots,10]$, the honeycomb diagrams display orthogonal transition lines confirming the individual control of each plunger to the corresponding QD, as shown in the left panel in Fig. 2(d).
@@
The goal of the plunger normalization layer is to ensure uniformity of charging voltages. This step gives three practical advantages. First, it allows the homogenization of the measurement window required for the next steps. Secondly, it enables to define all detuning ($\varepsilon_{i,j}$, with $i,j\in[1,\dots,10]$) and total-energy ($U_{i,j}$, with $i,j\in[1,\dots,10]$) axes of all double-QD pairs, often adopted for readout and initialization schemes, with the same simple $45^{\circ}$-rotation matrix. Thirdly, when a sizeable charge jump occurs, visualizing a charge state with a different charging voltage (due to the few-hole filling structure) immediately triggers the experimentalist or a surveying algorithm to perform a more in-depth verification of the charge state.
@@
Figure 4: Barrier fine virtualization (ON regime). (a) Sequence of CSDs N5 vs N8 as a function of J8 in the range [-110, 5] $\mathrm{mV}$ with respect to the dc reference point. (b) Sequence of CSDs N5 vs N8 as a function of K8 in the same range. The cyan points in (a) and (b) indicate the position of the interdots returned by the ML module and the magenta diamonds indicate the center of the tracked honeycomb. The shift in the (N8, N5) space of the central charge sector in panel (a) reveals imperfect virtualization over the large voltage range (115 $\mathrm{mV}$). Panel (b), where the finely virtualized gate K8 is adopted, shows a decreased susceptibility to the barrier gate voltage change. (c, d) Concise presentation of the evolution of the charge state while stepping J8 and K8, respectively. The finely calibrated K8 preserves the position of the charge state. (e) The fit to the center of the honeycomb positions for plunger gates N5 and N8 as a function of barrier gate J8 indicating a beyond-linear dependence for N5. (f, g) The quadratic coefficient $\sqrt{\alpha}$ and linear compensation coefficient $\beta$ respectively, with K8 vs. N5 and vs. N8 highlighted.
@@
The uniformity of charging voltages with a target constant spacing between consecutive transitions in a new virtual gate space spanned by the normalized plunger gates $\mathrm{N}_{i}$, with $i\in[1,\dots,10]$, is achieved through a diagonal virtualization matrix, $M_{3}$. The desired spacing across the charge transitions and interdots encompassing the central charge state of interest is set to $V_{D}=20$ $\mathrm{mV}$. The elements of matrix $M_{3}$ are determined using the slope extraction method described in Sec. II.4 applied to 2D CSDs acquired after the first plunger orthogonalization is completed. Since the uniformity of charging voltages is determined in an already orthogonalized plunger space $\mathrm{O}_{i}$, with $i\in[1,\dots,10]$, the resulting matrix $M_{3}$ is diagonal; see the right panel in Fig. 2(d) for $M^{-1}_{3}$. The homogenized spacing between charge transitions in the ($\mathrm{N}_{i}$, $\mathrm{N}_{j}$) space, with $i,j\in[1,\dots,10]$, is illustrated in the right panel of Fig. 2(e).
@@ III.3 Barrier Virtualization in the OFF-coupling regime
However, on occasion, the pixel classifier fails to identify interdots or identifies false positives [for examples of each, see the first and last 2D CSD in Fig. 3(a)]. Moreover, as the 2D CSD shifts due to crosstalk, interdots may shift into or out of the frame as the barrier voltage is changed. In such cases, the quality of the linear fit will be significantly affected, resulting in sub-optimal accuracy of the extracted slope. Thus, rather than finding the slopes directly, we rely on the local change in the interdots position, as described in Sec. II.5.

### v2

@@ II.3 Image processing
To extract from the experimentally acquired 2D CSDs information useful for automation, we use an ensemble of five convolutional neural network pixel classifiers [54] trained to distinguish each pixel as belonging to five different classes: horizontal, vertical, or diagonal transitions (corresponding to QD formed under left, right, or simultaneously under both plunger gates), interdots, and points where no transitions are present [45]. The pixel classifiers, trained using exclusively simulated data [55, 56], take as an input a small 2D plunger-plunger voltage scan obtained using a charge sensor, as shown in Fig. 1(c). It then labels each pixel within the scan with a 5D vector indicating the probability of belonging to each of the five possible charge transition classes.
Prior to running the pixel classifiers, the measured data are preprocessed. This involves a series of steps. First, the data are convolved with a Gaussian filter, and the gradient is taken. The resulting image is then mean normalized, and the normalized gradient is cropped to produce a $32\times 32$ pixel subimage.
Figure 2: Plunger gates orthogonalization and normalization. (a) Exemplary charge stability diagram spanned by the sensor-virtualized gates $\mathrm{vP}_{7}$ and $\mathrm{vP}_{3}$. In red, we overlay the output of the pixel classifier module for vertical transitions. The region circled in green indicates a single transition segment, while the blue dashed line indicates the composite line averaged over the three segments of the D7 transition line on the left-hand side of the image. The red line in the center of the image has been erroneously marked by the classifier. (b) The Hough transform of the ML model output. The individual transitions become thin bands, which overlap at a single point corresponding to the composite line (marked in blue). The three additional peaks (marked in green) correspond to the individual transitions on the left-hand side. Each pixel has a width of $0.17$ $\mathrm{mV}$ and a height of $0.35$ $\mathrm{mV}$. (c) The sum of the squares of the Hough transform, forming a bimodal distribution corresponding to the composite and individual transitions. (d) Inverse of the virtual matrix $M_{2}$ obtained from the Hough transforms (left) and a 2D CSD acquired with orthogonalized virtual plunger gates (right). (e) Inverse of the normalized virtual plunger matrix $M_{3}$ (left) and 2D CSD acquired with normalized virtual plunger gates. In the space spanned by $\mathrm{N}_{7}$ and $\mathrm{N}_{3}$, the charging voltages for the specific charge state are constant (20 $\mathrm{mV}$), allowing for uniform $x$ and $y$ axes.
@@
To identify the coordinates of the interdots, we implement a dynamic thresholding algorithm that selects the interdots based on the probabilities $\mathrm{P}(\ell)$ given by the ML module, where $\ell$ is the predicted class. The threshold $\delta$ is determined based on the mean $(\mu)$ and variance $(\sigma^{2})$ of a Gaussian fit to the log-transformed interdot class probability distribution,
@@
where $\eta$ is set dynamically in the range of $(2.5,4)$ in order to identify at least the expected number of interdots; see Fig. 1(f). In the end, a pixel is classified as belonging to an interdot if $\mathrm{P}(\ell)>\delta$. The final interdot coordinates in the 2D CSD are determined based on the center of mass of each cluster of points classified as an interdot, as illustrated in Fig. 1(g).
@@ II.4 Plunger virtualization
From a 2D CSD, the ML module identifies regions that correspond to horizontal and vertical transitions. For example, in Fig. 2(a), all pixels identified as belonging to vertical transitions are marked in red. This includes three left-hand side transitions and three right-hand side transitions, as well as an erroneously identified line in the center of the honeycomb. The ML output is then processed using the Hough transform, resulting in a representation where each transition becomes a thin band in the Hough space, with $\theta$ and $d$ corresponding to the angle and distance from the origin of the transition; see Fig. 2(b).
@@
The 2D CSDs often contain several nearly collinear transition lines that have slight offset relative to each other due to the presence of interdot transition, as shown in Fig. 2(a). This effect manifests itself visually in the Hough transformation as several small peaks, shown in Fig. 2(b) with green ovals, each corresponding to an individual transition. The large peak, shown in Fig. 2(b) with a blue oval, corresponds to a composite line that passes through the centers of the nearly collinear transitions [indicated as a dashed blue line in Fig. 2(a)]. The ability to distinguish between these two cases is essential, as the slope of the composite line will be different from the slopes of the individual transitions themselves, and this difference will be magnified as the size of the interdots increases.
When this occurs, $H_{\text{sq}}(\theta)$ will have a bimodal distribution, with the peak closest to $\theta=\pi/2$ ($\theta=0$) for the case of horizontal (vertical) transitions corresponding to the composite line, and the other peak reflecting the slope of the individual transition segments, as shown in Fig. 2 (c). Our implementation of MAViS is designed to prefer the peak corresponding to individual transitions so that the horizontal and vertical transitions are orthogonal in the final virtualized space, although in principle the composite line can be chosen instead. After finding the angle $\theta$ corresponding to the individual transitions, the slope of these transitions is then obtained via basic trigonometric functions. This process is repeated for all nearest-neighbor and next-nearest-neighbor plunger-plunger pairs, yielding the elements of the matrix $M_{2}^{-1}$ in Eq. (2).
@@ II.5 Barrier virtualization
The next layer of MAViS is the virtualization of the barrier gates with respect to the QD levels. Because of the barrier-to-QD capacitive couplings, nonvirtualized barriers result in a shift of the charge state within a CSD as the barrier gate is modified. Virtualized barriers are built to ensure that such a charge state, enclosed by four interdots, remains at the charge symmetry point despite varying the tunnel coupling between neighboring QDs.
Barrier virtualization involves determining the virtualization coefficients in Eqs. (4) and (5) that will correct for the shifts of the interdot positions as the barrier gates are adjusted. The same ML models trained on 2D CSDs obtained from plunger-plunger sweeps are leveraged for this task. First, the interdot locations in the 2D CSD are determined as described in Sec. II.3. To mitigate errors in tracking individual interdots, the center of the honeycomb formed by a set of four interdots is tracked instead. If the pixel classifier fails to identify one of the tracked interdots in one of the scans, coordinates for the corresponding interdot from the previous scan are used instead. Then, we derive the rate at which the charge state shifts as the barrier voltage is changed using one of two methods.
@@
However, local tracking is appropriate only for finding the linear fit coefficient. For the quadratic fit observed in the ON regime, a global tracking method is implemented. The global method involves direct tracking of the trajectories of the interdot transition points as the barrier gate is pulsed over multiple scans. The center coordinates as a function of the corresponding barrier voltages are then fit to a linear or quadratic curve, as desired. While missing or misclassified points can negatively affect the quality of the fit, virtualization in the OFF regime ensures less drastic shifts in the ON regime, making the fit less prone to errors. Moreover, with a limited number of points involved in the fit, a single misidentified point can substantially affect the resulting fit and lead to a miscalibration, which makes this approach less robust with noisy data. To overcome this limitation and ensure a high-quality fit, the virtualization in the ON regime is established based on a significantly denser sampling of 2D CSD for every pair of gates. Since only the neighboring gates need to be virtualized at this step, a larger number of measurements per gate pair can be afforded compared to the OFF regime virtualization.
@@ III Results
MAViS is designed to autonomously virtualize a set of sensors, plungers, and barrier gates used to define QDs. It operates through five distinct modules, each corresponding to a specific virtualization layer outlined in Table 1. At its core is an ML pixel classifier that extracts high-level feature representations from experimental data. MAViS leverages horizontal and vertical transition classes to determine the plunger-plunger virtualization coefficients, while the interdot class is used to identify and track interdot transitions for barrier virtualization. The performance of each virtualization layer is discussed in the remainder of this section, while we focus on the scalability and time requirements of MAViS in Supplemental Material [60].
@@ III.1 Charge sensors compensation
When searching for clear honeycomblike patterns in 2D CSDs measured via charge sensing, it is critical to maintain the rf-reflectometry charge sensors tuned to their maximum-sensitivity point, i.e., at the flank of one of their Coulomb peaks, throughout the whole voltage scan. The high sensitivity of charge sensing is ensured by calibrating the relative lever arm of each gate controlling the QD array to the charge sensors’ plunger gates. This step does not entail using ML algorithms and exploits only traditional analysis routines to track the charge sensors’ Coulomb peak position as a function of each barrier and plunger gate.
The charge sensor compensation matrix $M_{1}$ is determined in the first virtualization layer by obtaining the crosstalk of each gate to the four charge sensors. This is achieved by extracting the slope of the Coulomb peak position as a function of each gate. The slope of the linear shift in the $\mathbf{S}$ vs $\mathbf{P}$, and $\mathbf{S}$ vs $\mathbf{B}$ space is extracted and fed into the matrix $M^{-1}_{1}$, as shown in Supplemental Material [60]. As we move on to the next virtualization layers, only the sensor-virtualized $\mathbf{vP}$ and $\mathbf{vB}$ parameters are considered.
@@ III.2 Plunger gates virtualization
The plunger gate orthogonalization matrix $M_{2}$ is determined using the slope extraction method described in Sec. II.4. First, a 2D CSD acquired for each pair of plunger gates ($\mathrm{vP}_{i}$, $\mathrm{vP}_{j}$), with $i,j\in[1,\dots,10]$, and $i\neq j$, is processed by the ML module to detect the horizontal and vertical charge transition lines; see Fig. 2. Then, a Hough transform method is used to extract the slopes of all relevant charge transitions; see Fig. 2(b). The inverses of the slopes define the elements of the orthogonalization matrix; see the right-hand panel in Fig. 2(d). When scanning the newly defined virtual plunger gates ($\mathrm{O}_{i}$, $\mathrm{O}_{j}$), with $i,j\in[1,\dots,10]$, the honeycomb diagrams display orthogonal transition lines confirming the individual control of each plunger to the corresponding QD, as shown in the left-hand panel in Fig. 2(d).
@@
The goal of the plunger normalization layer is to ensure uniformity of charging voltages. This step gives three practical advantages. First, it allows the homogenization of the measurement window required for the next steps. Second, it enables us to define all detuning ($\varepsilon_{i,j}$, with $i,j\in[1,\dots,10]$) and total-energy ($U_{i,j}$, with $i,j\in[1,\dots,10]$) axes of all double-QD pairs, often adopted for readout and initialization schemes, with the same simple $45^{\circ}$-rotation matrix. Third, when a sizable charge jump occurs, visualizing a charge state with a different charging voltage (due to the few-hole filling structure) immediately triggers the experimentalist or a surveying algorithm to perform a more in-depth verification of the charge state.
@@
Figure 4: Barrier fine virtualization (ON regime). (a) Sequence of CSDs N5 vs N8 as a function of J8 in the range $[-110,5]$ $\mathrm{mV}$ with respect to the dc reference point. (b) Sequence of CSDs N5 vs N8 as a function of K8 in the same range. The cyan points in (a) and (b) indicate the position of the interdots returned by the ML module and the magenta diamonds indicate the center of the tracked honeycomb. The shift in the (N8, N5) space of the central charge sector in (a) reveals imperfect virtualization over the large voltage range (115 $\mathrm{mV}$). Panel (b), where the finely virtualized gate K8 is adopted, shows a decreased susceptibility to the barrier gate voltage change. (c), (d) Concise presentation of the evolution of the charge state while stepping J8 and K8, respectively. The finely calibrated K8 preserves the position of the charge state. (e) The fit to the center of the honeycomb positions for plunger gates N5 and N8 as a function of barrier gate J8 indicating a beyond-linear dependence for N5. (f), (g) The quadratic coefficient $\sqrt{\alpha}$ and linear compensation coefficient $\beta$, respectively, with K8 vs N5 and vs N8 highlighted.
@@
The uniformity of charging voltages with a target constant spacing between consecutive transitions in a new virtual gate space spanned by the normalized plunger gates $\mathrm{N}_{i}$, with $i\in[1,\dots,10]$, is achieved through a diagonal virtualization matrix $M_{3}$. The desired spacing across the charge transitions and interdots encompassing the central charge state of interest is set to $V_{D}=20$ $\mathrm{mV}$. The elements of matrix $M_{3}$ are determined using the slope extraction method described in Sec. II.4 applied to 2D CSDs acquired after the first plunger orthogonalization is completed. Since the uniformity of charging voltages is determined in an already orthogonalized plunger space $\mathrm{O}_{i}$, with $i\in[1,\dots,10]$, the resulting matrix $M_{3}$ is diagonal; see the right-hand panel in Fig. 2(d) for $M^{-1}_{3}$. The homogenized spacing between charge transitions in the ($\mathrm{N}_{i}$, $\mathrm{N}_{j}$) space, with $i,j\in[1,\dots,10]$, is illustrated in the right-hand panel of Fig. 2(e).
@@ III.3 Barrier virtualization in the OFF-coupling regime
However, on occasion, the pixel classifier fails to identify interdots or identifies false positives [for examples of each, see the first and last 2D CSD in Fig. 3(a)]. Moreover, as the 2D CSD shifts due to crosstalk, interdots may shift into or out of the frame as the barrier voltage is changed. In such cases, the quality of the linear fit will be significantly affected, resulting in suboptimal accuracy of the extracted slope. Thus, rather than finding the slopes directly, we rely on the local change in the interdots’ position, as described in Sec. II.5.